# PRIME: Scaffolding Manipulation Tasks with Behavior Primitives for Data-Efficient Imitation Learning

### v2

@@ Abstract
Imitation learning has shown great potential for enabling robots to acquire complex manipulation behaviors. However, these algorithms suffer from high sample complexity in long-horizon tasks, where compounding errors accumulate over the task horizons. We present PRimitive-based IMitation with data Efficiency), a behavior primitive-based framework designed for improving the data efficiency of imitation learning. caffolds robot tasks by decomposing task demonstrations into primitive sequences, followed by learning a high-level control policy to sequence primitives through imitation learning. Our experiments demonstrate that chieves a significant performance improvement in multi-stage manipulation tasks, with 10-34% higher success rates in simulation over state-of-the-art baselines and 20-48% on physical hardware. [^1] Additional materials are available at https://ut-austin-rpl.github.io/PRIME/.
@@
Imitation Learning, Deep Learning in Grasping and Manipulation, Deep Learning Methods
@@ I Introduction
Imitation learning (IL) has become a powerful paradigm for programming robots to perform manipulation tasks. Policies trained through imitation have exhibited diverse and complex behaviors, such as assembling parts [52], preparing coffee [54], making pizza [6], and folding cloth [49]. Deep IL methods aim at training policies that map sensory observations directly to low-level motor commands [4, 27, 6]. While conceptually simple, these methods usually require a large volume of human demonstrations [27], making them costly for tackling long-horizon tasks. Furthermore, the direct imitation of low-level motor actions leads to limited generalization abilities of the learned policy.
One solution to improve data efficiency and model generalization is incorporating temporal abstraction into policy learning [38]. Conventional methods afforded with temporal abstraction decouple learning new tasks into learning what subtasks to perform and how to achieve them. Among these methods, skills represent a popular form of temporal abstraction, providing a systematic approach to scaffolding complex tasks for robots. Skills serve as fundamental building blocks, capturing robot behaviors necessary for executing specific tasks, such as grasping an object. The first step of incorporating skills as temporal abstraction is to obtain a repertoire of motor skills that capture how to perform behaviors and serve as reusable building blocks applicable to a variety of robot tasks. With these motor skills, the problem of learning new tasks reduces to learning what behaviors to perform rather than how to perform them, which significantly reduces the complexity of the learning problem and facilitates generalization. Subsequently, the second step involves learning a policy for skill sequencing. To acquire skills, one popular approach is skill learning [30, 1, 35, 43], which learns low-level skills that capture short-horizon sequences of robot actions by learning either continuous latent skill representations [30, 1, 35] or a discrete set of skills with continuous parameters  [43, 55]. Prior work in skill learning extracts skills from a large amount of prior human data. While promising, a core limitation of these methods is the need for substantial human data to ensure the learned skills possess a high generalization capability.
@@
Figure 2: Method Overview. We develop a self-supervised data collection procedure that randomly executes sequences of behavior primitives in the environment. With the generated dataset, we train an inverse dynamics model (IDM) that maps an initial state $s$ and a final state $s^{\prime}$ from segments in task demonstrations to a primitive type $p$ and corresponding parameters $x$. To derive the optimal primitive sequences, we build a trajectory parser capable of parsing task demonstrations into primitive sequences using the learned IDM. Finally, we train the policy using parsed primitive sequences.
@@
Recent work has explored using robotic behavior primitives [31, 8, 5, 21] to scaffold manipulation tasks, such as movement primitives [14, 32], motion planning [47, 10, 20], and grasping systems [23, 2]. A behavior primitive is a parameterized module designed to capture a certain movement pattern, usually with explicit semantic meaning (e.g., grasping). The input parameters serve to instantiate the behavior primitive into a particular movement. These primitives enjoy the advantages of re-usability, modularity, and robustness toward variations. To utilize these primitives, recent work has proposed learning high-level policies that discover the optimal sequence of primitives using reinforcement learning (RL) [31, 7, 8]. However, RL requires expensive exploration even in the action space afforded by the primitives. Another notable line of work learns the policy with primitives from segmented demonstrations using imitation learning [45, 12, 24]. Segmented demonstrations require costly human efforts to manually segment demonstrations into primitive sequences.
@@
In response to the above challenges, we introduce PRIME (PRimitive-based IMitation with data Efficiency), a data-efficient imitation learning framework based on behavior primitives (see Fig. 1). We provide a small set of task demonstrations with raw sensory observations and a discrete collection of behavior primitives. Our framework consists of a two-step process for learning policies for new tasks: first, parsing task demonstrations as primitive sequences via a trajectory parser without the need for any human annotations, and subsequently, training a policy through imitation learning to predict the sequence of primitives (such as “push” and “grasp”) and their corresponding parameters given observations. By incorporating primitives, we break down long-horizon tasks into shorter sequences of primitives, significantly reducing the complexity and temporal horizon for imitation learning.
@@
In this work, PRIME does not require access to segmentation labels, rendering the parsing of demonstrations a challenging task. To segment demonstrations into primitive sequences, it is essential to establish a mapping from raw observation-action sequences to primitives. To generate the data necessary for learning the mapping, we introduce a self-supervised data collection procedure [37, 5] that randomly samples sequences of primitives to execute within the environment, effectively reducing the need for human efforts in collecting prior data. Subsequently, we train an Inverse Dynamics Model (IDM) [33, 53, 3, 34, 9] on the collected data, which maps pairs of states to primitives. We use this IDM with a dynamic programming algorithm to identify the optimal primitive sequences derived from task demonstrations.
@@
We evaluate our method’s effectiveness across long-horizon and contact-rich tabletop manipulation tasks, both in simulation and on real hardware. Our results highlight PRIME’s substantial performance gains over state-of-the-art imitation learning baselines in a low-data regime. In simulated tasks, we observe success rate increases ranging from 10.0% to 33.6%, and while on real robots, we see greater improvements, ranging from 20.0% to 48.3%. We further verify that our trajectory parser can effectively parse task demonstrations into primitive sequences that can be replayed to accomplish the task with success rates exceeding 90%. Moreover, our IDM generalizes to unseen environments, achieving performance levels comparable to those in training environments.
@@ II-A Learning from Demonstration
Learning from demonstration (LfD) has shown promise in robot manipulation tasks [40]. LfD aims to enable an agent to observe and replicate expert behavior to effectively achieve a designated task. Within the domain of LfD, a diverse and extensive range of approaches has emerged, encompassing imitation learning [27, 13, 25], demonstration-guided RL [35, 36, 46], and offline RL [1, 26, 17, 50]. These approaches that rely on demonstration guidance often necessitate a substantial number of expert demonstrations, thereby limiting their data efficiency. To reduce the burdens of collecting expert demonstrations, a common line of work learns from task-agnostic play data [22, 29, 30], which is more cost-effective to acquire but still demands a certain level of human supervision. Instead of relying on additional human data, we propose utilizing pre-defined primitives and data acquired from random primitive rollouts.
@@ II-B Skill-based Imitation Learning
Skill-based imitation learning extracts low-level temporally-extended sensorimotor behaviors as skills from expert demonstrations [51, 39] or task-agnostic play data [22, 29, 30] and emulates the high-level behavior observed in the expert demonstrations to guide the execution of these low-level skills. A common approach involves joint learning of low-level skills and a high-level policy [55, 30, 35], with skills acquired in an unsupervised manner [42, 41, 15, 16]. These skills can be either discrete [55] or continuous in a latent space [30, 35]. Unsupervised learning obviates the need for additional human annotation but often results in skills with limited reusability and low generalization capability. Alternatively, some research focuses on learning a high-level policy and low-level skills from structured demonstrations with additional segmentation labels using supervised learning, relying on either weak [45] or strong human supervision [12]. In our work, we use pre-built parameterized behavior primitives as low-level skills. These primitives are highly robust, reusable, and generalizable. Furthermore, our method requires only raw sensory demonstrations without the need for additional human annotations.
@@ II-C Learning with Behavior Primitives
One line of research focuses on policy learning with primitives, which involves augmenting the motor action space through the integration of parameterized primitives [31, 7, 11, 8, 19, 18]. Dalal et al. [8] proposes to manually specify a comprehensive library of robot action primitives. These primitives are carefully parameterized with arguments that are subsequently fine-tuned and learned by an RL policy. Similarly, Nasiriany et al. [31] augments standard RL algorithms by incorporating a pre-defined library of behavior primitives. Chitnis et al. [7] decomposes the learning process into learning a state-independent task schema, which represents a sequence of skills to execute, and a policy that selects the parameters of these skills based on the current state. In prior work, augmenting policy learning with parameterized primitives offers a promising approach to reducing the task horizon. However, the discrete-continuous augmented action space still imposes a significant exploration burden in RL. Recently, Chen et al. [5] proposes a model-based imitation learning framework that integrates primitives for solving stowing tasks. This approach only requires a single demonstration with a few key frames, but it involves a complex graph construction for Graph Neural Networks to predict forward dynamics, along with a specialized particle-based representation for object states. Another tangential work by Shi et al. [44] introduces a method for waypoint extraction in imitation learning. This approach decomposes demonstrations into sequences of waypoints, which are then interpolated through linear motion. The interpolated linear motion between waypoints can be regarded as a type of primitive. In contrast, our primitive-based framework can be viewed as a more versatile form of waypoint extraction, capable of encompassing a broader spectrum of skills.
@@ III Method
We introduce our primitive-based imitation learning framework, which decomposes complex, long-horizon tasks into sequences of primitives. We begin by formulating the problem and providing an overview of our framework. We then describe two components of our framework: the trajectory parser and the policy.
@@ III-A Problem Formulation
We formulate a robot manipulation task as a Parameterized Action Markov Decision Process (PAMDP) [28], defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},p_{0},\mathcal{R},\gamma)$ representing the continuous state space $\mathcal{S}$, the discrete-continuous parameterized action space $\mathcal{A}$, the transition probability $\mathcal{P}$, the initial state distribution $p_{0}$, the reward function $\mathcal{R}$, and the discount factor $\gamma$. In our setting, the motor action space is afforded by the primitives into discrete-continuous primitive action space, $a=(p,x),~{}a\in\mathcal{A},~{}p\in\mathcal{L},x\in\mathcal{X}_{p}$, where $\mathcal{L}$ is a discrete set of primitive types and $\mathcal{X}_{p}$ is the parameter space of primitive $p$. We aim to learn a policy, $\pi(a|s)=\pi(p,x|s)$, $s\in\mathcal{S}$, to maximize the expected sum of discounted rewards. We assume access to a small set of task demonstrations for imitation learning.
@@ III-B Trajectory Parser
We develop a trajectory parser to parse task demonstrations into primitive sequences. This parser comprises an IDM and a dynamic programming algorithm. The IDM learns the probability of mapping from segments of task demonstrations to primitives and the dynamic programming algorithm determines the optimal primitive sequence by maximizing the product of probabilities in the parsed primitive sequences.
@@ III-B1 Inverse Dynamics Model
To collect the training dataset for the IDM, we introduce a self-supervised data collection procedure by randomly executing primitives within the environment. Specifically, we deploy a random policy where, at each step, the policy uniformly samples a primitive type $p$ from the pre-built primitive set, randomly samples the parameters $x$ from the parameter space associated with $p$, and executes this instantiated primitive within the environment. This process yields a set of trajectories comprising the rollouts of these primitives. These primitive rollouts are subsequently utilized to train the IDM in a supervised manner. To gather data for the “other” category, we randomly sample a selection of state pairs from the generated trajectories.
Since we use a random policy for collecting each trajectory without any policy updating, the data collection process can be efficiently executed in parallel environments. Besides, in contrast to RL, which necessitates extensive exploration by the agent to accomplish the entire task, our data collection strategy merely requires the agent to interact effectively with objects, significantly reducing the exploration burden.
To enhance the quality of the training data, we filter out unsuccessful rollouts from the dataset using the success criteria of primitives. The success criterion for each primitive serves as a measure to assess the success of its execution. For example, a successful execution of a grasping primitive can be defined as the gripper securely holding an object upon completion. If we do not filter out failed rollouts, a significant portion of grasping rollouts within the dataset would involve grasping actions performed in the air, lacking effective interaction with objects. Failed rollouts would negatively impact the learned IDM, thereby degrading the accuracy and quality of the demonstration segmentation.
Moreover, gathering successful primitive rollouts through uniform parameter sampling can be prohibitively inefficient. To reduce the sampling burden and accelerate the data collection process, we incorporate a simple prior distribution consisting of a mixture of Gaussian distributions with the centers of the distributions aligning with the positions of task objects in the environment. The prior distribution directs the agent to interact more effectively with objects, improving the success rate of primitive rollouts. The pseudo-code for the data generation process is summarized in Alg. 1.
Algorithm 1 Self-Supervised Data Collection Procedure
@@
3:     $\mathcal{L}$: pre-built primitive set
4:     $\mathcal{C}_{p}$: number of episodes during data generation
5:     $\mathcal{M}$: horizon of episodes
6:     $\mathcal{K}$: number of negative samples for each episode
7:
8:  for $e\leftarrow 1,2,\cdots\mathcal{C}_{p}$ do
9:     for $i\leftarrow 1,2,\cdots\mathcal{M}$ do
10:        Sample a primitive type $p^{i}$ from $\mathcal{L}$
11:        Sample $x_{p}^{i}$ from parameter space of $p^{i}$
12:        Execute instantiated primitive $p^{i}(x_{p}^{i})$:
13:           Get $\tau^{i}=(s_{t_{i}},a_{t_{i}},s_{t_{i}+1},...,s_{t_{i+1}})$
14:        if is_success($\tau^{i},p^{i}$then
15:           $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t_{i}},s_{t_{i+1}},p^{i},x_{p}^{i})\}$
16:        end if
17:     end for
18:     Sequence all consecutive primitive rollouts $\{\tau^{i}\}_{i=1}^{M}$ into an episode trajectory:
19:        Get $\tau=(s_{0}=s_{t_{0}},a_{0},...,s_{t_{1}},a_{t_{1}},...,s_{t_{2}},...,s_{t_{M}})$
20:     for $k\leftarrow 1,2,\cdots K$ do
21:        Sample a segment $\tau^{\prime}=(s_{j},a_{j},...,s_{l})$ from $\tau$
22:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{j},s_{l},\texttt{other},\texttt{none})\}$
23:     end for
24:  end for

### v3

# IEEEabrv.bib

@@ Abstract
Imitation learning has shown great potential for enabling robots to acquire complex manipulation behaviors. However, these algorithms suffer from high sample complexity in long-horizon tasks, where compounding errors accumulate over the task horizons. We present PRIME (PRimitive-based IMitation with data Efficiency), a behavior primitive-based framework designed for improving the data efficiency of imitation learning. PRIME scaffolds robot tasks by decomposing task demonstrations into primitive sequences, followed by learning a high-level control policy to sequence primitives through imitation learning. Our experiments demonstrate that PRIME achieves a significant performance improvement in multi-stage manipulation tasks, with 10-34% higher success rates in simulation over state-of-the-art baselines and 20-48% on physical hardware. [^1] Additional materials are available at https://ut-austin-rpl.github.io/PRIME/.
@@
Imitation Learning, Deep Learning in Grasping and Manipulation, Deep Learning Methods.
@@ I Introduction
Imitation learning (IL) has become a powerful paradigm for programming robots to perform manipulation tasks. Policies trained through imitation have exhibited diverse and complex behaviors, such as assembling parts [zhao2023aloha], preparing coffee [zhu2022viola], making pizza [chi2023diffusionpolicy], and folding cloth [wang2023mimicplay]. Deep IL methods aim at training policies that map sensory observations directly to low-level motor commands [rt12022, mandlekar2021matters, chi2023diffusionpolicy]. While conceptually simple, these methods usually require a large volume of human demonstrations, making them costly for tackling long-horizon tasks. Furthermore, the direct imitation of low-level motor actions leads to limited generalization abilities of the learned policy.
One solution to improve data efficiency and model generalization is incorporating temporal abstraction into policy learning [precup2000temporal]. Conventional methods afforded with temporal abstraction decouple learning new tasks into learning what subtasks to perform and how to achieve them. Among these methods, skills represent a popular form of temporal abstraction, offering a systematic approach to decomposing complex tasks for robots. Skills serve as fundamental building blocks, capturing necessary robot behaviors for specific tasks, such as grasping an object. The first step of incorporating skills as temporal abstraction is to obtain a repertoire of motor skills that capture how to perform behaviors, serving as reusable building blocks for various tasks. This reduces the problem of learning new tasks to learning what behaviors to perform rather than how to perform them, simplifying the learning process and enhancing generalization. The second step involves learning a policy for skill sequencing. To acquire skills, one popular approach is skill learning, which learns low-level skills that capture short-horizon sequences of robot actions by learning either continuous latent skill representations [nasiriany2022learning, ajay2020opal, pertsch2021accelerating] or a discrete set of skills with continuous parameters  [shankar2020discovering, zhu2022bottom]. Prior work in skill learning extracts skills from a large amount of prior human data. While promising, a core limitation of these methods is the need for substantial human data to ensure the learned skills possess a high generalization capability.
@@
Figure 2: Method Overview. We develop a self-supervised data collection procedure that randomly executes sequences of behavior primitives in the environment. With the generated dataset, we train an IDM that maps an initial state $s$ and a final state $s^{\prime}$ from segments in task demonstrations to a primitive type $p$ and corresponding parameters $x$. To derive the optimal primitive sequences, we build a trajectory parser capable of parsing task demonstrations into primitive sequences using the learned IDM. Finally, we train the policy using parsed primitive sequences.
@@
Recent work has explored using robotic behavior primitives [nasiriany2022augmenting, dalal2021accelerating, chen2023predicting, luo2023multi] to decompose manipulation tasks, such as movement primitives [ijspeert2013dmps, neumann2014learning], motion planning [toussaint2015logic, garrett2021integrated, lozano2014constraint], and grasping systems [mahler2017dexnet2, bohg2013data]. A behavior primitive is a parameterized module designed to capture a certain movement pattern, usually with explicit semantic meaning (e.g., grasping). The input parameters instantiate the behavior primitive into a specific movement, with the output being a sequence of motor actions to control the robot. These primitives enjoy the advantages of re-usability, modularity, and robustness toward variations. To utilize these primitives, recent work has proposed learning high-level policies that discover the optimal sequence of primitives using reinforcement learning (RL) [nasiriany2022augmenting, chitnis2020efficient, dalal2021accelerating]. However, RL requires expensive exploration even in the action space afforded by the primitives and is unsafe to train on real robots. Another notable line of work learns the policy with primitives from segmented demonstrations using imitation learning [shiarlis2018taco, huang2019neural, mahmoudieh2020weakly]. Segmented demonstrations require costly human efforts to manually segment demonstrations into primitive sequences.
@@
In response to the above challenges, we introduce PRIME (PRimitive-based IMitation with data Efficiency), a data-efficient imitation learning framework based on behavior primitives (see Fig. 1). We provide a small set of task demonstrations with raw sensory observations and a discrete collection of behavior primitives. Our framework consists of a two-step learning process: first, parsing task demonstrations as primitive sequences via a trajectory parser without the need for any human annotations, and subsequently, training a policy through imitation learning to predict the sequence of primitives (such as “push” and “grasp”) and their corresponding parameters given observations. By incorporating primitives, we break down long-horizon tasks into shorter sequences of primitives, significantly reducing the complexity and temporal horizon for imitation learning.
@@
In this work, PRIME does not require access to segmentation labels, rendering the parsing of demonstrations a challenging task. To segment demonstrations into primitive sequences, it is essential to establish a mapping from raw observation-action sequences to primitives. To generate the data necessary for learning the mapping, we introduce a self-supervised data collection procedure [pinto2016supersizing, chen2023predicting] that randomly samples sequences of primitives to execute within the environment, effectively reducing the need for human efforts in collecting prior data. Subsequently, we train an Inverse Dynamics Model (IDM) [paster2020planning, zheng2023semi, brandfonbrener2023inverse, pavse2020ridm, du2024learning] on the collected data, which maps pairs of states to primitives. We use this IDM with a dynamic programming algorithm to identify the optimal primitive sequences derived from task demonstrations.
@@
We evaluate our method’s effectiveness in tabletop manipulation tasks, both in simulation and on real hardware. Our results highlight PRIME’s substantial performance gains over state-of-the-art imitation learning baselines in a low-data regime. In simulations, success rates increase by 10.0% to 33.6%, and on real robots, by 20.0% to 48.3%. We further verify that our trajectory parser can effectively parse task demonstrations into primitive sequences that can be replayed to accomplish the task with success rates exceeding 90%. Moreover, our IDM generalizes to unseen environments, achieving performance levels comparable to those in training environments.
@@ II-A Learning from Demonstration
Learning from Demonstration (LfD) has shown promise in robot manipulation tasks [ravichandar2020recent, chernova_gmm, paraschos2013probabilistic, traPPCA]. LfD aims to enable an agent to observe and replicate expert behavior to effectively achieve a designated task. Within the domain of LfD, a diverse and extensive range of approaches has emerged, encompassing imitation learning [mandlekar2021matters, hussein2017imitation, mandlekar2020learning], demonstration-guided RL [pertsch2021accelerating, pertsch2021guided, singh2020parrot], and offline RL [ajay2020opal, mandlekar2020iris, kumar2022pre, yu2022uds]. These approaches that rely on demonstration guidance often necessitate a substantial number of expert demonstrations, thereby limiting their data efficiency. To reduce the burdens of collecting expert demonstrations, a common line of work learns from task-agnostic play data [lynch2020learning, mees2022calvin, nasiriany2022learning], which is more cost-effective to acquire but still demands a certain level of human supervision. Instead of relying on additional human data, we propose utilizing pre-defined primitives and data acquired from random primitive rollouts.
@@ II-B Skill-based Imitation Learning
Skill-based imitation learning extracts low-level temporally-extended sensorimotor behaviors as skills from expert demonstrations [zhang2018deep, Rajeswaran-RSS-18] or task-agnostic play data [lynch2020learning, mees2022calvin, nasiriany2022learning] and emulates the high-level behavior observed in the expert demonstrations to guide the execution of these low-level skills. A common approach involves joint learning of low-level skills and a high-level policy, with skills acquired in an unsupervised manner [shankar2019discovering, shankar2020learning, konidaris2012robot, krishnan2017ddco]. These skills can be either discrete [zhu2022bottom] or continuous in a latent space [nasiriany2022learning, pertsch2021accelerating]. Unsupervised learning obviates the need for additional human annotation but often results in skills with limited reusability and low generalization capability. Alternatively, some research focuses on learning a high-level policy and low-level skills from structured demonstrations with additional segmentation labels using supervised learning, relying on either weak [shiarlis2018taco] or strong human supervision [huang2019neural]. In our work, we use pre-built parameterized behavior primitives as low-level skills, which are highly robust, reusable, and generalizable. Furthermore, our method requires only raw sensory demonstrations without the need for additional human annotations.
@@ II-C Learning with Behavior Primitives
One line of research focuses on policy learning with primitives, which involves augmenting the motor action space through the integration of parameterized primitives [nasiriany2022augmenting, chitnis2020efficient, hausknecht2015deep, dalal2021accelerating, lee2019learning, strudel2020learning]. Dalal et al. [dalal2021accelerating] propose to manually specify a comprehensive library of robot action primitives. These primitives are carefully parameterized with arguments that are subsequently fine-tuned and learned by an RL policy. Similarly, Nasiriany et al. [nasiriany2022augmenting] augments standard RL algorithms by incorporating a pre-defined library of behavior primitives. Chitnis et al. [chitnis2020efficient] decomposes the learning process into learning a state-independent task schema. The discrete-continuous augmented action space imposes a significant exploration burden in RL. Recently, Chen et al. [chen2023predicting] proposed an imitation learning framework that integrates primitives for solving stowing tasks, which involves a complex graph construction for Graph Neural Networks to predict forward dynamics. Another tangential work by Shi et al. [shi2023waypoint] decomposes demonstrations into sequences of waypoints, which are interpolated through linear motion. The interpolated linear motion between waypoints can be regarded as a type of primitive. In contrast, our primitive-based framework can be viewed as a more versatile form of waypoint extraction, capable of encompassing a broader spectrum of skills.
@@ III Method
We introduce PRIME, our primitive-based imitation learning framework, which decomposes complex, long-horizon tasks into concise, simple sequences of primitives. We begin by formulating the problem and providing an overview of our framework. We then describe two components of our framework: the trajectory parser and the policy.
@@ III-A Problem Formulation
We formulate a robot manipulation task as a Parameterized Action Markov Decision Process (PAMDP) [masson2016reinforcement], defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},p_{0},\mathcal{R},\gamma)$ representing the continuous state space $\mathcal{S}$, the discrete-continuous parameterized action space $\mathcal{A}$, the transition probability $\mathcal{P}$, the initial state distribution $p_{0}$, the reward function $\mathcal{R}$, and the discount factor $\gamma$. In our setting, the motor action space is afforded by the primitives into discrete-continuous primitive action space, $a=(p,x),~{}a\in\mathcal{A},~{}p\in\mathcal{L},x\in\mathcal{X}_{p}$, where $\mathcal{L}$ is a discrete set of primitive types and $\mathcal{X}_{p}$ is the parameter space of primitive $p$. We aim to learn a policy, $\pi(a|s)=\pi(p,x|s)$, $s\in\mathcal{S}$, to maximize the expected sum of discounted rewards. We assume access to a small set of task demonstrations for imitation learning.
@@ III-B Trajectory Parser
We develop a trajectory parser to parse task demonstrations into primitive sequences. This parser comprises an IDM and a dynamic programming algorithm. The IDM learns the probability of mapping from segments of task demonstrations to primitives. The dynamic programming algorithm determines the optimal primitive sequence by maximizing the product of probabilities in the parsed primitive sequences.
@@ III-B1 Inverse Dynamics Model
To collect the training dataset for the IDM, we introduce a self-supervised data collection procedure by randomly executing primitives and atomic motor actions within the environment. Specifically, our random policy either uniformly samples a primitive type $p$ from a pre-built set or selects a random atomic motor action, each with a 50% probability. If a primitive is sampled, its parameters $x$ are randomly chosen and the primitive is executed. Otherwise, the sampled motor action is executed directly. This process yields a set of trajectories comprising the rollouts of these primitives. These primitive rollouts are subsequently utilized to train the IDM in a supervised manner. To gather data for the “other” category, we randomly sample a selection of state pairs from the generated trajectories. Specifically, we uniformly sample $K$ state pairs in each episode and label them as “other” category, enabling us to balance the dataset by selecting an appropriate $K$ (see Alg. 1). To further address data imbalance during training, we reweight the data for each primitive type $p$ in the IDM training dataset by a factor of 1/(number of rollouts for primitive type $1/(\text{number of rollouts for primitive type }p)$.
The collected dataset is an offline dataset and differs from online RL samples. The data collection policy we used is not task-specific, generating a domain-specific dataset. Consequently, all tasks within the same domain share a single IDM, requiring only one dataset collection procedure.
To enhance training data quality, we filter out unsuccessful rollouts using the success criteria for each primitive. For example, a successful grasping primitive is defined by the gripper securely holding an object. Without this filtering, many rollouts would include ineffective actions, degrading the accuracy and quality of the learned IDM and demonstration segmentation.
Gathering successful primitive rollouts through uniform parameter sampling is inefficient. To reduce the sampling burden and speed up data collection, we use a prior distribution of Gaussian mixtures centered on task objects. This directs the agent to interact more effectively with objects, improving rollout success rates. The pseudo-code for the data generation process is summarized in Alg. 1.
@@
3:     $\mathcal{C}_{p}$: number of episodes during data generation
4:     $\mathcal{M}$: horizon of episodes
5:     $\mathcal{K}$: number of negative samples for each episode
6:
7:  for $e\leftarrow 1,2,\cdots\mathcal{C}_{p}$ do
8:     for $i\leftarrow 1,2,\cdots\mathcal{M}$ do
9:        Sample a primitive and its parameters $(p^{i},x_{p}^{i})$ or
an atomic motor action $a_{t_{i}}$ and set $p^{i}=\text{atomic}$
10:        Execute it and get $\tau^{i}=(s_{t_{i}},a_{t_{i}},s_{t_{i}+1},...,s_{t_{i+1}})$
11:        if $p^{i}\neq$ atomic and is_success($\tau^{i},p^{i}$then
12:           $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t_{i}},s_{t_{i+1}},p^{i},x_{p}^{i})\}$
13:        end if
14:     end for
15:     Sequence $\{\tau^{i}\}_{i=1}^{M}$ into an episode trajectory:
16:        Get $\tau=(s_{0}=s_{t_{0}},a_{0},...,s_{t_{1}},a_{t_{1}},...,s_{t_{2}},...,s_{t_{M}})$
17:     for $k\leftarrow 1,2,\cdots K$ do
18:        Sample a segment $\tau^{\prime}=(s_{j},a_{j},...,s_{l})$ from $\tau$
19:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{j},s_{l},\texttt{other},\texttt{none})\}$
20:     end for
21:  end for
Algorithm 1 Self-Supervised Data Collection Procedure